# AI-Powered Autonomous Weapons Risk Geopolitical Instability and Threaten AI Research

### v1

@@ 1 Introduction
With the rise of powerful generative AI models such as GPT-4 and Stable Diffusion, and continued progress in fields such as robotics and reinforcement learning, concerns have grown among both experts and the public about giving AI too much power. Academic concerns have thus far centered on threats in domains such as cybersecurity, biological weapons, disinformation, fraud, and hypothetical rogue artificial general intelligence (AGI) (Future of Life Institute, 2023; Bengio, 2023). Despite these general concerns, there has been relatively little attention given to specific recent developments from military and defense-industry groups, which have already begun to deploy next-generation AI-guided Autonomous Weapon Systems (AWS). Weapons falling under the AWS label have traditionally been either remotely operated (but AI-assisted) or only autonomous within a very narrow scope. However, new fully autonomous (unmanned) AWS in development remove or reduce the human role in the control and decision making process, with the goal of removing humans from the active battlefield en mass. For example, the Pentagon’s Replicator program for AI-based weapon “swarms” promises a drastic shift in warfare towards highly autonomous and cooperative AI units within the next few years (Greenwalt, 2023). Many of these new advanced AI and machine learning (ML) weapons systems are already seeing real-world deployment for the first time in the Ukraine War, and are being designed for every branch of the military and by many nations (Sharma, 2022; Greenwalt, 2023; Pfaff et al., 2023).
@@
To be clear, the ethical implications of such AWS capable of selecting and destroying targets without meaningful human control are considerable, and have been discussed by prior work at length (Scharre, 2018; Blanchard & Taddeo, 2022; Brown, 2023). Over the past 2-3 decades many voices have called for these systems to be banned entirely and/or governed by international treaties (Russell, 2023; Bode et al., 2023). We agree that the moral ideal would be that such systems were not developed and deployed, due to moral hazards in their use as well as the issues we highlight here. However, in this paper we focus on the scenario in which a middle ground is achieved: AWS are not banned despite moral reasons they ought to be (Geist, 2016), but manage to comply with basic military ethics and do not increase civilian casualties compared to conventional war (Scharre, 2018; Brose, 2020; Riesen, 2022). This assumption of no increased impact from war is not trivial to achieve—likely many AWS will fall short—but it provides the most conservative starting point for our discussion on the negative impacts of AWS. In 2024, major military powers have already committed to making autonomous weapon systems core components of their armed forces, and thus the time to discuss full AWS bans seems increasingly in the past (Greenwalt, 2023; Klare, 2023; Brose, 2020). The conversation now needs to center on responsible development and proliferation of the rapidly improving AWS systems that are already being developed and used by global militaries.
Official military statements and announced R&D projects make it clear that the direction of AWS development efforts both near- and long-term is the removal of human soldiers from direct combat roles, to reduce casualties and increase combat effectiveness (Greenwalt, 2023; Pfaff et al., 2023; Bachmann & Grant, 2023). While these goals are reasonable in isolation, a lack of public attention and transparency around the rapid and increasing pace of AWS development and employment risks humanity sleepwalking into a major AWS arms race between global powers. We will see these risks in the near future. China and Russia have given 2028-2030 as targets for major automatization of their militaries to begin, and while the USA is set to begin substantial deployment sooner, the country is planning for a long transformation timeline (Sharma, 2022; Greenwalt, 2023; Kania, 2020; Warren & Hillas, 2022).
@@
Dual-use AI technology Many AI algorithms are innately dual-use. Facial recognition, navigation, robotics, etc. Military interest in civilian research likely to grow. Improve university ethics oversight and transparency for military-funded AI research, and caution researchers against efforts to weaponize AI.
@@
Beyond the impacts AWS have on global stability, the importance of AWS for warfare will likely lead to major negative impacts on civilian AI research. AWS will become a revolutionary military technology, as did nuclear weapons, mechanized warfare, and others historically, but with important differences in the ease of AWS proliferation and impact on civilian technology development (Stern, 2023; Johnson, 2020; Laird, 2020; Wong et al., 2020). Recent work by Schneider argues that the prevailing military logic—that hegemonic power creates stability—can be highly erroneous when that power is based on technology that relies on a scarce or controlled resource, the capability/vulnerability paradox (Schneider, 2019). In the case of AWS, these key resources are AI experts and knowledge, access to relevant big data, and semiconductor manufacturing—all of which have historically been dominated by a handful of countries (Chu, 2023; Daniels, 2019; Khan et al., 2021; Farrell & Newman, 2022; Chahal et al., 2020). While these resources are becoming more available globally, governments with a current advantage are likely to try to secure that advantage by preventing their spread. We are likely to see a rise in export restrictions, publication oversight and redaction, and knowledge compartmentalization in the field of AI as a result (Bengio, 2023; Rickli & Ienca, 2021; Warren & Hillas, 2020).
@@
Longer term, we may see AI resources treated analogously to how oil resources have fueled conflict since the advent of industrialized warfare (Schneider, 2019). Indeed, we see hints in current events that this misguided idea is already starting to spread (Templeman & Mastro, 2023; Klingler-Vidra, 2023; Triolo, 2023), and some militaries are already seeking to use civilian AI experts to speed development of domestic AWS (Pfaff et al., 2023). China for example has declared Military-Civil Fusion as foundational to updating its military with AWS capabilities, and major Chinese universities have been roped into AI weapons programs (Konaev et al., 2023; Can & Vieira, 2022; Warren & Hillas, 2021; Fedasiuk, 2021).
@@
To reduce and prevent these outcomes, action by AI researchers, policymakers, and the public will be needed. Below we propose several policies and actions to take in Section 4 and Table 1, and briefly review highlights here. First, the best way to avoid AWS-driven conflicts is for nations to avoid developing or fielding AWS that operate independently rather than in cooperation with humans, especially in large scale or prolonged conflicts. Second, the international community needs to build consensus around AWS autonomy levels, and which levels are acceptable or unacceptable and in what circumstances. Third, there is an urgent need for transparency around AWS capabilities and levels of human oversight in actual use—“Human-in-the-loop” is becoming a meaningless phrase that arms manufacturers and militaries use to reassure the public. Today, “human-in-the-loop” risks shifting from implying hands-on remote control to automated targeting where a human presses a “go” button with minimal insight into what the AWS is doing or why (Bode & Watts, 2023; Bode et al., 2023; Boys, 2023; Abraham, 2024). Fourth, universities and companies worldwide must draw lines and declare publicly what level of military-civilian overlap and collaboration on AWS-related research is acceptable (D’Agostino, 2024).
@@
The remaining structure of this article is as follows: Section 2 discusses the state of AWS development and deployment today, as well as likely near-future directions. Section 3 discusses why limiting AWS proliferation is difficult, and the ineffectiveness of restricting access to cutting-edge hardware and ML expertise to prevent it. Section 4 discusses several alternative approaches to mitigate the risk posed by advanced AWS short of a total ban.
@@ 2.1 Airborne AWS
Since then, the number of LMs has rapidly expanded, with at least 24 models of varied size and sophistication across 16 countries, and the capabilities of the AI systems used by these systems have increased considerably (Bode & Watts, 2023). Most concerning, the range of targets for LMs has expanded from radar systems to armored vehicles and even specific human personnel that can be identified and targeted by deep learning and facial recognition without operator involvement. The Kargu-2 developed by Turkey is one of the foremost systems with this capability, and was deployed on battlefields in Libya (Longpre et al., 2022; Bode & Watts, 2023). Other notable models such as the Rafael Spike Firefly and the IAI Rotem L are optimized for targeting in urban environments, and at least 14 other models have specifications suggesting they can autonomously target and follow or execute individual human targets (Bode & Watts, 2023). Both Ukraine and Russia have deployed LMs of various levels of sophistication and autonomy in their current war in large numbers (Bode & Watts, 2023; Cotovio et al., 2024). Notably, the former commander in chief of the Ukrainian military recently highlighted LMs and anti-LM defenses such as jamming as critical to their war effort (Zaluzhnyi, 2023). He also highlighted the Russian Lancet LM, which uses a western-built Nvidia Jetson TX2 ML compute module to autonomously track targets, as being particularly difficult to defend against (Zaluzhnyi, 2023; Faragasso, 2023). A major issue in assessing LMs is that governments and manufacturers are often vague about the capabilities of LM AI-targeting systems, and whether autonomous modes without human oversight have been used operationally (Bode & Watts, 2023; Hambling, 2023a; Sayler, 2021; Longpre et al., 2022).
@@
One example system is the XQ-58A Valkyrie developed by Kratos Defense and Security Solutions to be a “loyal wingman” to manned aircraft, defending them and performing offensive actions that would otherwise risk the manned aircraft, first demonstrated in 2019 (Wang et al., 2020; Klare, 2020, 2023; Newdick & Rogoway, 2023). Notably, Kratos has claimed it can produce several hundred Valkyrie units per year at per-unit costs between \$2-5 million, cheaper than manned aircraft, most traditional drones like the American Reaper UAV, and even some missiles (Jordan, 2021; Lyu & Zhan, 2022; Boys, 2023). These units can scout, are equipped with weapons for defensive fire, have an operational range of  3000 miles, and can even deploy smaller unmanned aircraft for strike or low-level reconnaissance (Newdick & Rogoway, 2023). Other major candidates are even cheaper, and are intended to be deployed in large numbers using autonomous swarming capabilities, such as Shield AI’s V-Bat and Boeing’s MQ-28 (Gruszczak & Kaempf, 2023; Saylam & Gündoğdu, 2023; Bachmann & Grant, 2023; Sud, 2020; Castillo, 2022). In addition, unmanned versions of the F-16 have been demonstrated recently, including the first dogfight with a human in 2024 and with deployment estimated for 2025 (DARPA, 2023; Brown, 2023; Kohler, 2024). US allies are developing similar systems, such as the Blue Bear Ghost line of autonomous UAVs (BlueBear, 2018).
@@ 2.2 Ground AWS
The USA and other NATO nations have driven most recent UGV developments of note (Sharma, 2022). In 2017, the US Army declared its mid-term (2021–30) priorities include the need to achieve the ability to perform fully automated convoy operations supported by unmanned combat vehicles and support vehicles that can navigate difficult terrain and carry advanced payloads (U.S. Army Training and Doctrine Command, 2017). Significant progress has been made towards this goal. The most notable recent unmanned ground vehicle (UGV), already deployed on the Ukrainian battlefield (Milrem Robotics, 2023a), is the Milrem THeMIS unit capable of autonomous targeting and ordinance deployment, as well as other support and sensing tasks (Mathiassen et al., 2021; Milrem Robotics, 2023b). Milrem has sought to make the unit more autonomous for faster threat response and to enable group coordination in small autonomous swarms.
Several other UGV AWS are in late development stages across the US, Europe and India, with major demonstrations or deployments planned (Sharma, 2022). For example, the MAARS Grey Muzzle autonomous weaponized robot developed by QinetiQ that is equipped with machine guns and grenade launchers for direct fire support roles. Currently, the platform features an onboard AI system for autonomous navigation and waypoint following, but the weapons require human input before firing (QinetiQ, 2023). Legged quadruped UGVs also have shown great promise for use in more complex terrain and environments (Hambling, 2023b; Gilead, 2023). Lastly, Pegasus Intelligence is developing a portfolio of diverse AWS, some of which are amphibious or can transition between flight and ground modes (Pegasus Intelligence, 2023). Looking forward, recent progress in subfields of ML and robotics, such as autonomous vehicle research, has greatly improved UGV capabilities and spurred additional development, with several recent DARPA programs testing self-driving abilities in military technology (Bisht, 2022).
@@ 3 AWS Proliferation and Threats to Academic Research
Presently, most forms of AWS are deployed by few countries, and only in limited numbers. As such, their military significance is relatively low—AWS capabilities are not core to any nation’s national defense at present. We anticipate this to change over time, possibly within 5-10 years, with AWS taking a more central role in many armed forces. Further, we expect AWS to proliferate widely, and perform key roles in many armies around the world (Scharre, 2018). This importance means that the technologies that drive AWS, machine learning and robotics, become critical defense technologies. As such, this defense value will influence AI research worldwide, and raises the prospect of national security restrictions on AI researchers. In this section, we argue why advanced AWS, once developed, will proliferate widely, and why policies that aim to regulate AI research to prevent AWS proliferation will be both futile and deleterious to the field.
@@
Firstly, regarding proliferation, the technologies needed (for example) to allow a tracked UGV such as the TheMIS to maneuver, detect, and fire on enemy troops with a usefully-high success rate already exist as an extension of methods developed for autonomous driving and computer vision applications (Mathiassen et al., 2021). Likewise, UAVs detecting targets and directing strikes against them without human intervention are already tractable- indeed, multiple loitering munitions advertise some degree of autonomous targeting functionality as described in Section 2. While non-trivial integration work remains to implement such capabilities more generally, the hardware and software engineering talent required is not scarce or restricted for countries with large educated workforces.
@@
Secondly, ML software proliferates easily and widely and is cheap to develop relative to many military procurement programs (Warren & Hillas, 2020; Bengio, 2023). The key breakthroughs enabling advanced AWS now rather than in past decades come from the field of machine learning, where major advances in algorithms have enabled new robotic applications such as warehouse logistics and remote survey and monitoring UAVs and UGVs (Schedl et al., 2021; del Cerro et al., 2021). In the near future, academic and industry researchers aim to expand that frontier to include human-assistant robots capable of navigating complex indoor environments (Fan et al., 2020), dexterous object manipulation (Andrychowicz et al., 2020; Nagabandi et al., 2020), and even surgery (Saeidi et al., 2022), with much effort going into developing such systems. These are worthy applications, but many of these algorithms are inherently “dual-use” in that they can be applied to AWS development by militaries and defense contractors (Hoijtink & Planqué-van Hardeveld, 2022; Gómez de Ágreda, 2020). While it is a good thing that most of the ML software ecosystem remains open source and freely available, this does mean that the core software needed for AWS development is also readily available.
@@
One major form these misplaced efforts to restrict knowledge take is immigration and research visa restrictions on civilian scientists. Indeed, such visa bans have already been put in place for some Chinese researchers by the USA for example (Hawkins, 2024). Such bans are not just ineffective for the reasons given above, but are pointlessly cruel to the people who are targeted, and can backfire and increase proliferation rather than reduce it. As a historical example, Qian Xuesen was a Chinese-American rocket scientist who was unjustly expelled from the USA during the early Cold War when rocket science suddenly became a field of critical national security interest, and who as a result became the father of the Chinese rocketry program under coercion after his deportation from the USA (BBC News, 2020).
@@
More broadly applicable, however, is the fact that AWS, like most robotics currently, do not require high end compute capacity. While much attention has accrued to LLMs such as ChatGPT which run on large clusters of expensive and scarce cutting-edge GPUs, most robotics and other edge ML research needs only modest hardware (Chen & Ran, 2019). These algorithms and the chips they run on are most relevant to AWS development, because both civilian and military robotics must reckon with limited volume, mass, power, and time budgets for on-robot compute. The chips suitable for these applications must be both small and low power, running models much smaller that state of the art LLMs, and will overlap heavily with those used for common civilian applications.
As a concrete example, Shield AI, who produce the cutting-edge V-Bat AWS platform capable of autonomous swarming behavior, lists the on-platform compute requirements for autonomy as follows: “AI Performance: 100 TOPS, GPU: 1024-core Nvidia Ampere architecture GPU + 32 Tensor Cores, CPU: 8-core Arm® Cortex®-A78AE v8.2 64-bit CPU 2MB L2 + 4MB L3” (Shield AI, 2023). These specifications correspond exactly to Nvidia’s Jetson Orin NX embedded edge compute module, which debuted in 2023 on Samsung’s 8nm process from 2018, and is representative of the current standard for embedded GPU compute. The performance of this module falls far below the most recent performance-based export restrictions on GPUs the US government has imposed, with the threshold for export control at 4800 TOPS, over 48 times more powerful than the Jetson Orin NX’s 100 TOPS (Department of Commerce, 2023). By comparison, Chinese semiconductor manufacturer SMIC advertises a 7nm-class process, while multiple Chinese silicon design companies have GPU accelerator chips in production or announced, making the production of a Chinese domestic chip roughly equivalent to the Jetson Orin NX likely in the near future.

### v2

@@ 1 Introduction
With the rise of powerful generative AI models such as GPT-4 and Stable Diffusion, and continued progress in fields such as robotics and reinforcement learning, concerns have grown among both experts and the public about giving AI too much power. Academic concerns have thus far centered on threats in domains such as cybersecurity, biological weapons, disinformation, fraud, and hypothetical rogue artificial general intelligence (AGI) (Future of Life Institute, 2023; Bengio, 2023). Despite these general concerns, there has been relatively little attention given to specific recent developments from military and defense-industry groups, which have already begun to deploy next-generation AI-guided Autonomous Weapon Systems (AWS). Weapons falling under the AWS label have traditionally been either remotely operated (but AI-assisted) or only autonomous within a very narrow scope. However, new fully autonomous (unmanned) AWS in development remove or reduce the human role in the control and decision making process, with the goal of removing humans from the active battlefield en mass. For example, the Pentagon’s Replicator program for AI-based weapon “swarms” promises a drastic shift in warfare towards highly autonomous and cooperative AI units within the next few years (Greenwalt, 2023). AWS can serve many battlefield roles, although human-targeting lethal AWS (LAWS) have received most public attention (Taddeo & Blanchard, 2022). Many of these new advanced AI and machine learning (ML) weapons systems are already seeing real-world deployment for the first time in the Ukraine War, and are being designed for every branch of the military and by many nations (Sharma, 2022; Greenwalt, 2023; Pfaff et al., 2023).
@@
The ethical implications of such AWS capable of selecting and destroying targets without meaningful human control are considerable, and have been discussed by prior work at length (Scharre, 2018; Blanchard & Taddeo, 2022; Brown, 2023). Over the past 2-3 decades many voices have called for these systems to be banned entirely and/or governed by international treaties (Russell, 2023; Bode et al., 2023). We agree that the ideal would be that such systems were not developed, due to moral hazards in their use as well as the issues we highlight here. However, in this paper we focus on the scenario in which a middle ground is achieved: AWS are not banned despite moral reasons they ought to be (Geist, 2016), but manage to comply with basic military ethics and do not increase civilian casualties compared to conventional war (Scharre, 2018; Brose, 2020; Riesen, 2022). This assumption that AWS will not increase civilian casualties is an optimistic one —likely many AWS will fall short of that goal— but it provides the most conservative starting point for our discussion on the indirect negative impacts of AWS. In 2024, major military powers have already committed to making autonomous weapon systems core components of their armed forces, and thus the time to discuss full AWS bans seems increasingly in the past (Greenwalt, 2023; Klare, 2023; Brose, 2020). The conversation now needs to center on responsible development and proliferation of the rapidly improving AWS systems that are already being developed and used by global militaries.
Official statements and announced R&D projects make it clear that the direction of AWS development efforts both short- and long-term is the removal of human soldiers from direct combat roles, to reduce casualties and increase combat effectiveness (Greenwalt, 2023; Pfaff et al., 2023; Bachmann & Grant, 2023). While these goals are reasonable in isolation, a lack of public attention and transparency around the rapid and increasing pace of AWS development and employment risks humanity sleepwalking into an AWS arms race between global powers. We will see these risks in the near future. China and Russia have given 2028-2030 as targets for major automatization of their militaries to begin, and while the USA is set to begin deployment sooner, the country is planning for a long-term transformation (Sharma, 2022; Greenwalt, 2023; Kania, 2020; Warren & Hillas, 2022).
@@
Dual-use AI technology Many AI algorithms are innately dual-use. Facial recognition, navigation, robotics, etc. Military interest in civilian research is likely to grow. Improve university ethics oversight and transparency for military-funded AI research, and caution researchers against efforts to weaponize AI.
@@
Beyond the impacts AWS have on global stability, the importance of AWS for warfare will likely lead to major negative impacts on civilian AI research. AWS will become a revolutionary military technology, as did nuclear weapons, mechanized warfare, and others historically, but with important differences in the ease of AWS proliferation and impact on civilian technology development (Stern, 2023; Johnson, 2020; Laird, 2020; Wong et al., 2020). Recent work by Schneider argues that the prevailing military logic—that hegemonic power creates stability—can be highly erroneous when that power is based on technology that relies on a scarce or controlled resource, the capability/vulnerability paradox (Schneider, 2019). In the case of AWS, these key resources are AI experts and knowledge, access to relevant big data, and semiconductor manufacturing—all of which have historically been dominated by a handful of countries (Chu, 2023; Daniels, 2019; Khan et al., 2021; Farrell & Newman, 2022; Chahal et al., 2020). While these resources are becoming more available globally, governments with a current advantage are likely to try to secure that advantage by preventing their spread. We may see a rise in export restrictions, publication oversight and redaction, and knowledge compartmentalization in the field of AI as a result (Bengio, 2023; Rickli & Ienca, 2021; Warren & Hillas, 2020).
@@
Longer term, we could see AI resources treated analogously to how oil resources have fueled conflict since the advent of industrialized warfare (Schneider, 2019). Indeed, we see hints in current events that this misguided idea is spreading (Templeman & Mastro, 2023; Klingler-Vidra, 2023; Triolo, 2023), and some militaries are already seeking to use civilian AI experts to speed development of domestic AWS (Pfaff et al., 2023). China for example has declared Military-Civil Fusion as foundational to updating its military with AWS capabilities, and major Chinese universities have been roped into AI weapons programs (Konaev et al., 2023; Can & Vieira, 2022; Warren & Hillas, 2021; Fedasiuk, 2021).
@@
To reduce and prevent these outcomes, action by AI researchers, policymakers, and the public will be needed. Below we propose several policies and actions to take in Section 4 and Table 1, and briefly review highlights here. First, the best way to avoid AWS-driven conflicts is for nations to avoid developing or fielding AWS that operate independently rather than in cooperation with humans, especially in large scale or prolonged conflicts. Second, the international community needs to build consensus around AWS autonomy levels, and which levels are acceptable or unacceptable and in what circumstances. Third, there is an urgent need for transparency around AWS capabilities and levels of human oversight in actual use—“Human-in-the-loop” is becoming a meaningless phrase that arms manufacturers and militaries use to reassure the public. Today, “human-in-the-loop” is shifting from implying hands-on remote control to automated targeting where a human presses a “go” button with minimal insight into what the AWS is doing or why (Bode & Watts, 2023; Bode et al., 2023; Boys, 2023; Abraham, 2024). Fourth, universities and companies worldwide must draw lines and declare publicly what level of military-civilian overlap and collaboration on AWS-related research is acceptable (D’Agostino, 2024).
@@
The structure of this article is as follows: Section 2 discusses the state of AWS development and deployment today, as well as near-future directions. Section 3 discusses why limiting AWS proliferation is difficult, and the ineffectiveness of restricting access to ML hardware and expertise to prevent it. Section 4 discusses several approaches to mitigate the risk posed by advanced AWS short of a total ban.
@@ 2.1 Airborne AWS
Since then, the number of LMs has rapidly expanded, with at least 24 models of varied size and sophistication across 16 countries, and the capabilities of the AI systems used by LMs have increased considerably (Bode & Watts, 2023). Most concerning, the range of targets for LMs has expanded from radar systems to armored vehicles and even specific human personnel that can be identified and targeted by facial recognition without operator involvement. The Kargu-2 developed by Turkey is one of the foremost systems with this capability, and was deployed on battlefields in Libya (Longpre et al., 2022; Bode & Watts, 2023). Other notable models such as the Rafael Spike Firefly and the IAI Rotem L are optimized for targeting in urban environments, and at least 14 other models have specifications suggesting they can autonomously target and follow or execute individual human targets (Bode & Watts, 2023). Both Ukraine and Russia have deployed LMs of various levels of sophistication and autonomy in their current war in large numbers (Bode & Watts, 2023; Cotovio et al., 2024). Notably, the former commander in chief of the Ukrainian military recently highlighted LMs and anti-LM defenses such as jamming as critical to their war effort (Zaluzhnyi, 2023). He also highlighted the Russian Lancet LM, which uses a western-built Nvidia Jetson TX2 ML compute module to autonomously track targets, as being particularly difficult to defend against (Zaluzhnyi, 2023; Faragasso, 2023). A major issue in assessing LMs is that governments and manufacturers are often vague about the capabilities of LM AI-targeting systems, and whether autonomous modes without human oversight have been used operationally (Bode & Watts, 2023; Hambling, 2023a; Sayler, 2021; Longpre et al., 2022).
@@
One example system is the XQ-58A Valkyrie developed by Kratos Defense and Security Solutions to be a “loyal wingman” to manned aircraft, defending them and performing offensive actions that would otherwise risk the manned aircraft, first demonstrated in 2019 (Wang et al., 2020; Klare, 2020, 2023; Newdick & Rogoway, 2023). Notably, Kratos has claimed it can produce several hundred Valkyrie units per year at per-unit costs between \$2-5 million, cheaper than manned aircraft, most traditional drones like the American Reaper UAV, and even some missiles (Jordan, 2021; Lyu & Zhan, 2022; Boys, 2023). These units can scout, are equipped with weapons for defensive fire, have an operational range of  3000 miles, and can even deploy smaller unmanned aircraft for strike or low-level reconnaissance (Newdick & Rogoway, 2023). Other major candidates are even cheaper, and are intended to be deployed in large numbers using autonomous swarming capabilities, such as Shield AI’s V-Bat and Boeing’s MQ-28 (Gruszczak & Kaempf, 2023; Saylam & Gündoğdu, 2023; Bachmann & Grant, 2023; Sud, 2020; Castillo, 2022). In addition, unmanned versions of the F-16 have been demonstrated recently, with deployment estimated for 2025 (DARPA, 2023; Brown, 2023; Kohler, 2024). US allies are developing similar systems, such as the Blue Bear Ghost line of autonomous UAVs (BlueBear, 2018).
@@ 2.2 Ground AWS
The USA and other NATO nations have driven most recent UGV developments of note (Sharma, 2022). In 2017, the US Army declared its mid-term (2021–30) priorities include achieving the ability to perform fully automated convoy operations using unmanned combat and support vehicles that can navigate difficult terrain and carry advanced payloads (U.S. Army Training and Doctrine Command, 2017). The most notable recent unmanned ground vehicle (UGV), already deployed on the Ukrainian battlefield (Milrem Robotics, 2023a), is the Milrem THeMIS unit, which is capable of autonomous targeting and ordinance deployment as well as other support and sensing tasks (Mathiassen et al., 2021; Milrem Robotics, 2023b). Milrem has sought to make the unit more autonomous for faster threat response and to enable coordination in autonomous swarms.
Several other UGV AWS are in late development stages across the US, Europe and India, with major demonstrations or deployments planned (Sharma, 2022). For example, the MAARS Grey Muzzle autonomous weaponized robot developed by QinetiQ that is equipped with machine guns and grenade launchers for direct fire support roles. This system features an onboard AI system for autonomous navigation and waypoint following, but the weapons require human input before firing (QinetiQ, 2023). Legged quadruped UGVs also have shown great promise for use in more complex terrain and environments (Hambling, 2023b; Gilead, 2023). Lastly, Pegasus Intelligence is developing a portfolio of diverse AWS, some of which are amphibious or can transition between flight and ground modes (Pegasus Intelligence, 2023). Looking forward, recent progress in subfields of ML and robotics, such as autonomous vehicle research, has greatly improved UGV capabilities and spurred additional development, with several recent DARPA programs testing self-driving abilities in military technology (Bisht, 2022).
@@ 3 AWS Proliferation and Threats to Academic Research
Presently, most forms of AWS are deployed by few countries, and only in limited numbers. As such, their military significance is relatively low—AWS capabilities are not core to any nation’s national defense at present. We anticipate this will change over time, possibly within 5-10 years, with AWS taking a more central role in how war is waged. Further, we expect AWS to proliferate widely, and perform key roles in many armies around the world (Scharre, 2018). If AWS are core to national defense, this means the technologies that power AWS capabilities, machine learning and robotics, are critical defense technologies. This means that developments in those fields have major military implications, which raises the prospect of national security restrictions on AI research and researchers worldwide. In this section, we argue why advanced AWS, once developed, will proliferate widely, and why policies that aim to regulate AI research to prevent AWS proliferation will be both futile and deleterious to the field.
@@
Firstly, regarding proliferation, the technologies needed (for example) to allow a tracked UGV such as the TheMIS to maneuver, detect, and fire on enemy troops with a usefully-high success rate already exist as an extension of methods developed for autonomous driving and computer vision applications (Mathiassen et al., 2021). Likewise, UAVs detecting targets and directing strikes against them without human intervention is already tractable- indeed, multiple loitering munitions advertise some degree of autonomous targeting functionality as described in Section 2. While non-trivial integration work is needed to implement such capabilities for any given AWS, the hardware and software engineering talent required is not scarce or restricted, as evidenced by the multitude of AI and ML products being developed by companies worldwide.
@@
Secondly, ML software proliferates easily and widely and is cheap to develop relative to many military procurement programs (Warren & Hillas, 2020; Bengio, 2023). The key breakthroughs enabling advanced AWS now rather than in past decades come from the field of machine learning, where major advances in algorithms have enabled new robotic applications such as warehouse logistics and remote survey and monitoring UAVs and UGVs (Schedl et al., 2021; del Cerro et al., 2021). In the near future, academic and industry researchers aim to expand that frontier to include human-assistant robots capable of navigating complex indoor environments (Fan et al., 2020), dexterous object manipulation (Andrychowicz et al., 2020; Nagabandi et al., 2020), and even surgery (Saeidi et al., 2022), with much effort going into developing such systems. These are worthy applications, but many of these algorithms are inherently “dual-use” in that they can be applied to AWS development by defense contractors (Hoijtink & Planqué-van Hardeveld, 2022; Gómez de Ágreda, 2020). While it is a good thing that most of the ML software ecosystem remains open source and freely available, this does mean that the core software needed for AWS development is also readily available.
@@
One major form these misplaced efforts to restrict knowledge take is immigration and research visa restrictions on civilian scientists. Indeed, such visa bans have already been put in place for some Chinese researchers by the USA (Hawkins, 2024). Such bans are not just ineffective for the reasons given above, but are pointlessly cruel to the people who are targeted, and can backfire and increase proliferation rather than reduce it. As a historical example, Qian Xuesen was a Chinese-American rocket scientist who was unjustly expelled from the USA during the early Cold War when rocket science suddenly became a field of critical national security interest, and who as a result became the father of the Chinese rocketry program under coercion after his deportation from the USA (BBC News, 2020).
@@
However AWS development, like most robotics currently, does not require high end compute capacity. While much attention has accrued to LLMs such as ChatGPT which run on large clusters of expensive and scarce cutting-edge GPUs, most robotics and other edge ML research needs only modest hardware (Chen & Ran, 2019). These algorithms and the chips they run on are most relevant to AWS development, because both civilian and military robotics must reckon with limited volume, mass, power, and time budgets for on-robot compute. The chips suitable for these applications must be both small and low power, running models much smaller than state of the art LLMs, and will overlap heavily with those used for common civilian applications.
As a concrete example, Shield AI, who produce the cutting-edge V-Bat AWS platform capable of autonomous swarming behavior, lists the on-platform compute requirements for autonomy as follows: “AI Performance: 100 TOPS, GPU: 1024-core Nvidia Ampere architecture GPU + 32 Tensor Cores, CPU: 8-core Arm® Cortex®-A78AE v8.2 64-bit CPU 2MB L2 + 4MB L3” (Shield AI, 2023). These specifications correspond to Nvidia’s Jetson Orin NX edge compute module, which debuted in 2023 on Samsung’s 8nm process from 2018, and is representative of the current standard for embedded GPU compute. The performance of this module falls far below the most recent performance-based export restrictions on GPUs the US government has imposed, with the threshold for export control at 4800 TOPS, over 48 times more powerful than the Jetson Orin NX’s 100 TOPS (Department of Commerce, 2023). By comparison, Chinese semiconductor manufacturer SMIC advertises a 7nm-class process, and multiple Chinese silicon design companies have GPU accelerator chips in production or announced, making the production of a Chinese domestic chip equivalent to the Jetson Orin NX likely in the near future.